Fermi National Laboratory
 

Operation Summary for February 7, 2011 - February 9, 2011

 

Monday Feb. 7

The day shift began with Operations monitoring the Tevatron (TEV) store 8482 (153.34E30) and the Recycler stash (262.6E10), with the Antiproton Source (Pbar) stacking (17.0E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8482 at 2:40 PM, put the TEV into HEP shot setup, and established store 8483 at 3:45 PM with an average initial luminosity (AIL) of 341E30. An operating system update issue interrupted MiniBooNE beam operations for 3 hours.

During the evening shift, a TEV quadrupole power supply (QDD1) developed a fuse fault. Operators prepared Switchyard enclosure G2 for access.

Tuesday Feb. 8

The midnight shift began with Operations monitoring store 8483 (121.27E30) and the stash (235.7E10), with Pbar stacking (29.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8483 at 7:05 AM. TEV personnel conducted abort kicker testing after operators had decelerated the TeV to 150 GeV. At the end of the kicker tests, Operations put the TEV into shot setup.

During the day shift, Operations established store 8484 at 8:34 AM with an AIL of 338E30. Interlock technicians conducted safety system tests in Meson.

During the evening shift, Operators prepared M01 for access. Operations reported that a MI LCW pump at MI-50 tripped off.

Wednesday Feb. 9

The midnight shift began with Operations monitoring store 8484 and the stash (408.0E10), with Pbar stacking (19.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8484 at 12:05 AM and put the TEV into shot setup, but a Linac RF station (LRF1) tripped off. Linac personnel came in and investigated the LRF1 failure; they reported that a power surge inside the modulator forced the replacement of several components. Operations reported that a MI-8 line quadrupole power supply (Q847) tripped off; an EE Support technician came in and replaced the controller chassis. Operations reported that the Linac beam returned at 7:10 AM. Operators disabled stacking and prepared AP-0 for an access into the water cage. Operations put the TEV into shot setup at 7:45 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct slipstacking and gap clearing studies and he will allow safety personnel to conduct Booster beam detector tests. On Thursday, 2/10/11, the Run Coordinator will allow MI an access to make RF repairs in the tunnel. Other machines may have an opportunity to make necessary repairs during this time.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/09/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory